Consider the following equation: Mg(s)+2H₂O(l)→Mg(OH)₂(s)+H₂(g). Find the volume of H₂ produced at STP by the reaction of 76.4 g of magnesium. *
1
Expert's answer
2020-11-17T10:36:16-0500

At Standard Temperature and Pressure (STP), 1 mole of any gas will occupy a volume of 22.4 L (Vm).


Solution:

The balanced chemical equation:

Mg(s) + 2H2O(l) → Mg(OH)2(s) + H2(g)

According to the equation: n(Mg) = n(H2)


Moles of Mg = n(Mg) = Mass Mg / Molar mass of Mg

The molar mass of Mg is 24.305 g mol-1

Thus,

n(Mg) = m(Mg) / M(Mg) = 76.4 g / 24.305 g mol-1 = 3.1434 mol

n(Mg) = 3.1434 mol


Hence,

n(H2) = n(Mg) = 3.1434 mol


Moles of H2 = n(H2) = Volume of H2 / Vm

Volume of H2 = n(H2) × Vm = 3.1434 mol × 22.4 L mol-1 = 70.412 L = 70.4 L

V(H2) = 70.4 L


Answer: 70.4 L of H2